Startup Night 17 : UT 20140915

Mostly clear at sunset, RHhigh and variable. We were open briefly but had to close.

LUCI1 Calibrations

luci.20140915.NNNN.fits

Darks to check readnoise...
  • 0001 0002 1x 4.0s O2DCR ==> 12.60 e-
  • 0003 0004 1x12.0s MER10 ==> 8.43 e-

Poor seeing

04:50 Running LBCB shutter tests with lights on in dome. ~5 X 10 exposures. First set had an RPC error. The instrument seemed to reconnect, but when we tried to run the OB again it indeed crashed. Once we recovered we got through the rest of the images with no problems and no shutter failures.

05:37 Still closed for humidity.

06:20 Starting to dry up. Opening soon. Seeing is now good!

LUCI Checks

Before updating the transform
  • 0007 1x4.0s blind
  • 0008 1x10.0s blind
  • ==> Use 2x10.0s with Br_gam filter
0009 0028 dice5_4x using old transform

Condensation blob on LUCI1 window is growing at 68% RH

0029 2x10.0s blind for dark

0030 0054 dice5_5x with new transform

LUCI1 -> blind mask, N375, mirror, both blind filters.

08:42 Switch to LBC

Scripts in /home/LBTO/SciPrep/LBCIQfields
  • run RA23h30m_Dec51d00m.ob to copoint after DOFPIA
  • run ra2330d51.ob once to cycle through filters
09:01 Run RA23h30m_Dec51d00m.ob but abort, just to point the telescope

09:04 DOFPIA, /first, /X2

Red has too much coma, need to correct by eye...+2500 nm Z8 (coma)
See logs in ~LBCeng/FPIAlogs/*2014-09-15.Log

09:25 co-pointed

09:26 ra2330d51.ob

Take ROTCEN data
  • 10:00 DOFPIA, /X2
  • rotatorcenter_pa0 R:100703 B:100706
  • rotatorcenter_pa180 R:100906 B:100907
Take a couple sets of superfoc (0 180 90 270) Calib_OBs/SUPERFOC
  • 10:10 DOFPIA, /X2
  • 10:15 RB_rV_superfoc7.ob
  • 10:21 RB_rV_superfoc7_pa180.ob
  • 10:29 DOFPIA
  • 10:34 RB_rV_superfoc7_pa90.ob
  • 10:41 RB_rV_superfoc7_pa270.ob
10:51 DOFPIA

Scripts in /home/LBTO/SciPrep/LBCIQfields
  • 10:56 run RA03h20m_Dec47d00m.ob to copoint
  • 11:01 ra0320d47_sbv_rIi.ob
  • 11:18 DOFPIA
  • 11:23 ra0320d47_Ugr_Vrz.ob
  • 11:39 DOFPIA
  • 11:45 ra0320d47_sVB_FYr (reduced to 1 dither pos for ea. filter)
Take some superfoc data here
  • 11:51 DOFPIA
  • 11:56 RB_rV_superfoc7.ob
  • 12:01 RB_rV_superfoc7_pa180.ob
12:08 Done! Closing dome.

Analysis

Olga's analysis of the superfoc data is at:

http://ohia.as.arizona.edu/~kuhn/SuperfocRes/2014Bsuperfoc.pdf

Christian's plots of the superfoc data are at:

http://abell.as.arizona.edu/~lbtsci/cv/lbciq20140915sf.html

-- %USERSIG{DavidThompson - 2014-09-15}%

Comments

 
Topic revision: r5 - 01 Oct 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback